find the compute integral of p2x+q2y=z;using charpits method


1
Expert's answer
2021-02-25T01:51:58-0500

Let's solve the problem

"f = p^2x+q^2y-z = 0"

"\large\frac{dp}{-p+q^2} = \large\frac{dq}{-q+q^2}=\large\frac{dz}{-2(p^2x+q^2y)}= \large\frac{dx}{-2px} = \large\frac{dy}{-2q}"

"\large\frac{2pxdp+p^2xdx}{2px(-p+p^2)+p^2(-2px)}= \large\frac{2qydq+q^2dy}{2qy(-q+q^2)+q^2(-2qy)}"

"\large\frac{d(p^2x)}{-2p^2x} = \large\frac{2(q^2y)}{-2qy}"

"\large\frac{}{}" "\large\frac{d(p^2x)}{p^2x} = \large\frac{d(q^2y)}{q^2y}"

"log(p^2x)=log(q^2)+loga"

"p^2x = q^2ya"

"aq^2y+q^2y = z \space\space \space\space \space\space \space\space q^2y(1+a)=z"

"q = \large[\large\frac{z}{(1+a)y}]^{\frac{1}{2}}"

"p^2 = \large\frac{q^2ya}{x}"

"p = q\large(\frac{ya}{x})^\frac{1}{2} = [\frac{2a}{(1+a)x}]^{\frac{1}{2}}"

"dz = pdx+qdy"

"dz = \large\{\frac{za}{(1+a)x}\}" "dx + \large\{\frac{za}{(1+a)y}\}" "dy"

Answer: "(1+a)^{1/2}\sqrt{z} = \sqrt{a}\sqrt{x}+\sqrt{y} + b"